OpenStudy (anonymous):

The area of a circle is 35 in². The length of the radius is _____in. 2.5 in 2.7 in 3.1 in 3.3 in

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

(35/pi)^(1/2)

OpenStudy (anonymous):

@pooja195 hello?

OpenStudy (anonymous):

so divide by pi then take the square root

OpenStudy (anonymous):

so i get 11.1464968152

OpenStudy (anonymous):

? hello

OpenStudy (anonymous):

\[\sqrt{\frac{ 35 }{ \pi }}=3.33\]

OpenStudy (anonymous):

so the last option

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Please everyone help me with another problem need help with 2 or 3 more

pooja195 (pooja195):

sryy but @ⓐ❦ is right